OODR SPECTROSCOPY OF c̃^{1}A_{1} CH_{2}

ZHONG WANG, GREGORY E. HALL, \underline{TREVOR J. SEARS}^{a}, Chemistry Department, Brookhaven National Laboratory, Upton, NY 11973-5000; and KAORI KOBAYASHI, Department of Physics, University of Toyama, 3190 Gofuku, Toyama 930-8555, Japan.

Optical-Optical Double Resonance spectroscopy has been used to search for rovibrational levels in the c̃^{1}A_{1} state of CH_{2} which are lifetime broadened due to predissociation. In the scheme used, an assigned rovibronic transition in the b̃^{1}B_{1} − ã^{1}A_{1} system was monitored using frequency-modulated laser transient absorption, while a pulsed ns OPO laser was scanned through visible wavelengths. The electric field of the pulsed laser caused transient shifts and broadening, due to the ac Stark effect in the upper level of the probed transition, when it was tuned to be resonant with an allowed upward transition connecting to the probed level. The sum of the two optical frequencies was chosen to match the approximately known bond dissociation energy of ã^{1}A_{1} CH_{2}. However, nearly all of the double resonance transitions detected involved a different mechanism: transient depletion of population in the lower level of the probed transition. These give information on previously undetected, J -selected, levels in the b̃-state. Analysis of the data provides energies for higher rovibronic levels in the b̃-state. Reasons for the non-observation of ac Stark resonances will also be discussed.

aWork at Brookhaven National Laboratory was carried out under Contract No. DE-AC02-98CH10886 with the U.S. Department of Energy and supported by its Division of Chemical Sciences, Office of Basic Energy Sciences.